What is the pH of a solution containing 0.50 M ammonia and 0.50 M ammonium chloride? What will the pH be after 25mL of 0.20 M of hydrochloric acid has been added to 100.0mL of the solution containing the ammonia and ammonium chloride? What will the pH be after 18mL of 0.25 M of lithium hydroxide has been added to 100.0mL of the solution containing the ammonia and ammonium chloride? The ionization constant for ammonia is 1.8x10-5.
Answer #1

@ NH3 | Nity formo a basse buffer POH= pkb (NH3) + loa [nity t] < conjugate world I og Ints ] < base = 4.75 + log 005070-50 =

fing) Malarity of added LPOH = (18 ro1_) 60.25M). 1. (10000+ 18) mL = 0.038 M Nitz-H + L+ oH > Nity + H2O + pt concentration

Hope you find it helpful